# Spectrally Deconfounded Random Forests

### v2

@@ 4.3 Spectrally Deconfounded Random Forests
The next natural step is to utilize spectrally deconfounded regression trees (SDTree) to construct spectrally deconfounded Random Forests (SDForests) for estimating arbitrary functions. Random Forests were introduced by Breiman (2001) and have been successfully employed in numerous applications. The idea is to combine multiple regression trees into an ensemble to decrease variance and obtain a smoother function. Ensembles work best if the different models are independent of each other. To decorrelate the regression trees as much as possible from each other, we have two mechanisms. The first one is bagging (Breiman 1996), where we train each regression tree on an independent bootstrap sample of the observations, i.e., we draw a random sample of size $n$ with replacement from the observations. The second mechanic to decrease the correlation is that only a random subset of the covariates is available for each split. Before each split, we sample $\mathsf{mtry}\leq p$ from all the covariates and choose the one that reduces the loss the most from those. Since we already have Algorithm 1, it only takes minor changes to build an SDForest with it. In Algorithm 1 before line 13, we sample a set $p_{\mathsf{mtry}}$ of size $\mathsf{mtry}$ from the covariates and in line 13, we only check in this set of randomly sampled covariates for the best split, e.g., $j\in p_{\mathsf{mtry}}$. This procedure gives the spectrally deconfounded Random Forest in Algorithm 2. We predict with all the trees separately and use the mean over all trees as the Random Forest prediction. This gives the estimated function
@@
     $SDtree_{t}\leftarrow$ SDTree from Algorithm 1 with random set of covariates of size $\mathsf{mtry}$ at each split using $X^{t}$
@@ 5 Empirical Results
$f^{0}(X)\coloneqq\sum_{j=1}^{p}\mathds{1}_{\{j\in\mathcal{J}_{s}\}}\sum_{k=1}^{K}a_{j,k}\cos(0.2k\cdot x_{j})+b_{j,k}\sin(0.2k\cdot x_{j})$ (13)
where $\mathcal{J}_{s}$ is a random subset of ${1,\ldots,p}$ of size four being the parents of $Y$ (among the $X$ variables). We simulate with $n=1000$, $p=500$, and $q=20$. The entries of $\mathbf{E}\in\mathbb{R}^{n\times p}$, $\mathbf{H}\in\mathbb{R}^{n\times q}$, $\delta\in\mathbb{R}^{q}$, and $\Gamma\in\mathbb{R}^{q\times p}$ are sampled i.i.d. from a Gaussian with expectation zero and $\sigma=1$. The additional noise $\mathbf{\nu}\in\mathbb{R}^{n}$ is sampled i.i.d. from a Gaussian with mean zero and $\sigma_{\nu}=0.1$. For the four causal parents, we sample the coefficients $a_{j,k}$ and $b_{j,k}$ uniformly on $[-1,1]$ with the number of basis functions fixed at $K=2$ for the additive function.
@@
In Figure 3, we compare the variable importance of the SDForest and the classical Random Forest for one simulation run. The variable importance of a covariate is calculated as the sum of the loss decrease resulting from all splits that divide a region using that covariate. The mean of the variable importance of all trees results in the variable importance for the forest. For the SDForest, we report the reduction of the spectral loss (3) instead of the MSE. For the SDForest, the four most important covariates are also the four true causal parents of $Y$. Three of those also have a clear separation from the remaining 497 covariates. For the classical Random Forest, the variable importance for the true causal parents lies within the bulk of spurious covariates. The most important covariate among the true causal parents is only on rank 102 and we would have to use almost all the covariates to include all the causal parents in the selected set.
@@
In addition to screening for the sparse causal set among a large number of covariates, we can also look at the functional dependence of the response $Y$ on the causal parents. We use partial dependence plots (Friedman 2001) to visualize the partial dependence. The idea is to predict $\hat{f}(\mathbf{X})$ for each observation and vary $\mathbf{X}_{j}$ on an interval while we keep the other covariates at the observed values. This gives for every observation a different function of $X_{j}$. The mean over all the observations can then be shown as a representative marginal effect of $X_{j}$ on the response $Y$. Figure 5 shows how the SDForest estimates the true causal function. Especially for covariate 34, the estimated function approximates the true function well. Covariate 108, the covariate with only slightly higher importance than the bulk of spurious covariates, has almost a constant influence on $Y$. This shows some limitations of estimating a sparse causal relationship in the presence of hidden confounding and high dimensionality. Such a nuanced true function as for variable 108 might be hard to estimate with the additional disturbance of confounding and noise. The classical Random Forest distributes the estimated function among all the spurious covariates and estimates almost a constant function for all four causal parents. The partial dependence plots for the estimated classical Random Forest are shown in Figure 6.
@@
Figure 6: Partial dependence plots as in Figure 5 but now with the estimated classical Random Forest by ranger for the four true causal parents of the response $Y$. The classical Random Forest essentially leads to constant functions, whereas the true dashed lines are varying.
5.1 Performance of SDForests
To quantify the performance of the SDForests in estimating the causal function, we perform simulation studies varying the different dimensions in the simulation. The default dimensions are $n=500$, $p=500$, and $q=20$, and we randomly draw data according to model (13). Each of those dimensions is varied separately to estimate the dependence of the performance on these different factors. Every experiment is repeated 200 times, where every time, the whole data-generating process is newly drawn at random. The performance is measured by the mean distance of the true causal function and the estimated function evaluated at 500 test observations $f_{mse}\coloneqq\frac{1}{n_{test}}\sum_{i=1}^{n_{test}}(f^{0}(x_{test,i})-\hat{f}(x_{test,i}))^{2}$. The performance of the SDForests and the classical Random Forests depending on different dimensions is shown in Figure 7. Theorem 2 shows how the spectral objective (3) behaves as the dimensionality $p$, and the number of observations $n$ grows to infinity. The term $R_{n}$ goes asymptotically with $\min(\sqrt{n},\sqrt{p})$ to zero. So, we need both $n$ and $p$ to grow for consistency. With this simulation study, we look at the dependence of the performance on those quantities in practice. Figure 7a shows the error distribution depending on the sample size. The SDForests perform clearly better in estimating the causal function, and the error decreases with a larger sample size. We see the dependence of the performance on the number of covariates in Figure 7b. At $p>20$, we see how the SDForests start to perform significantly better than the classical Random Forests. When increasing $p$ above 50, the error no longer seems to decrease, which hints that there is a certain threshold of $p$ after which deconfounding is successful. Even with only the four causal parents as covariates, the SDForests do not perform worse than the classical versions, and we really do not seem to lose anything with SDForests, even in low-dimensional settings. With stronger confounding, estimating the causal function becomes increasingly difficult. Not only does the bias increase but so does the variance in the response. Therefore, we expect the error to increase with increasing confounding even for the SDForests. Figure 7c shows this behavior with a simulation study, where we increase the number of confounding variables. For $q=0$, the setting corresponds to the classical model without confounding, and both the classical Random Forests and the SDForests perform really well. It is important to note that we do not lose much, even when the data is not confounded, by applying the spectral deconfounding. However, as can be seen for increasing confounding, we can gain a lot if there is hidden confounding present. In Figure 7d, we follow up on the assumption of dense confounding. Here, instead of having random effects of the confounders on all the covariates, we simulate the data with only a random subsample of the covariates being affected by confounding. The denser the confounding becomes, the better SDForests perform. With around 200 covariates affected (40%), the SDForests start performing similarly as if there were no confounding, while with sparser confounding, SDForest still outperforms ranger.
@@ 6 Single-Cell Data
We apply SDForest and ranger to Single-cell gene expression data to compare the resulting model on the scRNA-seq dataset for the cell RPE1 generated by Replogle et al. (2022). We use the preprocessed and filtered dataset provided by Chevalley et al. (2025). As the response, we choose the gene for EIF1 following the arguments of Shen et al. (2023) conjecturing that EIF1 might be a leaf node in the causal graph of genes. Using the observational data without any interventional gene knockouts ($n=11485$ observations), we use all the other gene expressions ($p=382$) as predictors and fit both ranger and SDForest. For both methods, we fit 500 trees using default parameters. To test for robustness against dense confounding, we synthetically perturb the original cleaned and filtered dataset. For this, we construct $\mathbf{X}_{\tau}\coloneqq\mathbf{X}+\mathbf{H}\Gamma\tau$ and $\mathbf{Y}_{\tau}\coloneqq Y+\mathbf{H}\delta\tau$, where $\mathbf{Y}$ is the original expression of EIF1 and $\mathbf{X}$ are the other original gene expressions used for prediction. The entries of $\mathbf{H}\in\mathbb{R}^{n}$, $\Gamma\in\mathbb{R}^{1\times 382}$, and $\delta\in\mathbb{R}$ are sampled i.i.d. from $\mathcal{N}(0,1)$ while we vary $\tau$ to increase the added dense confounding. We fit both methods to $\mathbf{X}_{\tau}$ and $\mathbf{Y}_{\tau}$ with increasing $\tau$ and analyze how much the estimated functions change. The change in the estimated functions is calculated as the change of the out-of-bag predictions of $f^{\text{\emph{ranger}}}_{\tau}(\mathbf{X}^{\tau})$ and $f^{\text{SDF}}_{\tau}(\mathbf{X}^{\tau})$ depending on $\tau$
@@
We repeat the synthetic confounding of the data by sampling 20 times $\mathbf{H}$, $\Gamma$, and $\delta$ and increasing $\tau$. The distributions of the function changes are shown in Figure 8. With no added confounding, the difference between the two functions is small (see caption of Figure 8). This hints at a finding that there is no major dense confounding present. As we increasingly add dense confounding, we clearly see how the estimated function by ranger changes, while SDForest demonstrates robustness against this added perturbation.
@@ 7 Conclusion
We propose the Spectrally Deconfounded Random Forest algorithm SDForest with R-package SDModels (Ulmer & Scheidegger 2025) to estimate direct regression functions in high-dimensional data in the presence of dense hidden confounding. This can be used to screen for relevant covariates among a large set of variables, and the procedure provides robustness against unobserved confounding, gaining much in the confounded case but losing little if no confounding is present. SDModels provides functions such as regularization paths and stability selection to screen for relevant covariates as well as partial dependence plots to understand their relation to the variable of interest. We demonstrate the empirical behavior of SDForests in various settings, including challenging cases with sparse confounding, to illustrate some fundamental limitations (which are not found to be clearly worse than those of classical Random Forests).
Many potentially confounded high-dimensional applications are about classification instead of regression. Currently, SDForest is only applicable for regression tasks, and it is important to extend this methodology to classification. Another open question is whether one can combine spectral deconfounding with Quantile Regression Forests (Meinshausen 2006) or Distributional Random Forests (Hothorn & Zeileis 2021, Ćevid et al. 2022) to gain access to prediction intervals or constructing confidence intervals using techniques as in Guo et al. (2022) for the linear case and Näf et al. (2023) using Random Forests.
@@ Acknowledgment
We thank Mathieu Chevalley for his support in applying our method to the single-cell dataset.
@@ Appendix A Approximation of splitting criteria
In the $M$th iteration of Algorithm 1, we have to find in each region the split that reduces the loss the most. This means we have $M$ regions in each iteration, each with an optimal split to choose from. When we split the region with the optimal split, we have two new regions. Now, due to the spectral transformation, the samples, as well as the different regions containing a subset of the samples, are not independent. Therefore, to truly find the next best split, we would need to determine the optimal split and its corresponding loss decrease in each region anew. This is done in Figure 9 using the method SDT2. In our studies and as the default in the R-package, we only estimate the optimal split and its loss decrease for the two new regions and reuse all the other estimates from the previous iteration. This saves substantial computational time and we argue that a previously good split stays reasonable. We compare the performance of the two options in Figure 9. We simulate data following the confounding model (1) and with the same parameter as in Section 5 but using a random regression tree for $f^{0}$ as in Equation 8. The random regression tree is grown using random splits until there are ten leaf nodes. While the reestimation of all splits in SDT2 might be a bit better, we do not see a significant decrease in performance when using the computationally much more efficient approximation SDT1. In the case of smooth underlying regression functions, we expect to see similar behavior (because we can consider the approximation error of a smooth function with a tree that remains unaffected and the estimation error of the best-approximated tree function, which is analogous to the discussion above).

### v3

@@ 4.3 Spectrally Deconfounded Random Forests
The next natural step is to utilize spectrally deconfounded regression trees (SDTree) to construct spectrally deconfounded Random Forests (SDForests) for estimating arbitrary functions. Random forests have been introduced by Breiman (2001) and have been successfully employed in numerous applications. The idea is to combine multiple regression trees into an ensemble to decrease variance and obtain a smoother function.
Ensembles work best if the different models are independent of each other. To decorrelate the regression trees as much as possible, we employ two mechanisms. The first one is bagging (Breiman 1996), where we train each regression tree on an independent bootstrap sample of the observations, i.e., we draw a random sample of size $n$ with replacement from the observations. The second mechanism to decrease the correlation is that only a random subset of the covariates is available for each split. Before each split, we sample $\mathsf{mtry}\leq p$ from all the covariates and choose the one that reduces the loss the most from those.
It only takes minor changes to Algorithm 1 to build an SDForest. In Algorithm 1 before line 13, we sample a set $p_{\mathsf{mtry}}$ of size $\mathsf{mtry}$ from the covariates, and in line 13, we only check in this set of randomly sampled covariates for the best split, e.g., $j\in p_{\mathsf{mtry}}$. This procedure yields the spectrally deconfounded Random Forest, as outlined in Algorithm 2. We predict with all the trees separately and use the mean over all trees as the Random Forest prediction, resulting in
@@
  Find $Q^{t}$ using $X^{t}$
  $SDtree_{t}\leftarrow$ SDTree from Algorithm 1 with random set of covariates of size $\mathsf{mtry}$ at each split using $X^{t}$ and $Q^{t}$.
@@ 5 Empirical Results
In this section, we analyze with a simulation study how well SDForests estimate a true causal function in comparison to classical Random Forests. Subsection 5.1 shows qualitatively how SDForest can screen for and estimate a sparse causal effect. The ability to estimate the causal function with respect to different dimensions and on the dense confounding assumption is examined quantitatively in subsection 5.2.
@@
$f^{0}(X)\coloneqq\sum_{j=1}^{p}\mathds{1}_{\{j\in\mathcal{J}_{s}\}}\sum_{k=1}^{K}(a_{j,k}\cos(0.2k\cdot x_{j})+b_{j,k}\sin(0.2k\cdot x_{j}))$ (13)
where $\mathcal{J}_{s}$ is a random subset of $1,\ldots,p$} of size four being the parents of $Y$ (among the $X$ variables). We simulate with $n=1000$, $p=500$, and $q=20$. The entries of $\mathbf{E}\in\mathbb{R}^{n\times p}$, $\mathbf{H}\in\mathbb{R}^{n\times q}$, $\delta\in\mathbb{R}^{q}$, and $\Gamma\in\mathbb{R}^{q\times p}$, in the confounding model (1), are sampled i.i.d. from a Gaussian with expectation zero and $\sigma=1$. The additional noise $\mathbf{\nu}\in\mathbb{R}^{n}$ is sampled i.i.d. from a Gaussian with mean zero and $\sigma_{\nu}=0.1$. For the four causal parents, we sample the coefficients $a_{j,k}$ and $b_{j,k}$ uniformly on [1,1][-1,1] with the number of basis functions fixed at $K=2$ for the additive function.
@@
5.1 Qualitative Results
In Figure 3, we compare the variable importance of the SDForest and the classical Random Forest for one simulation run. The variable importance of a covariate is calculated as the sum of the loss decrease resulting from all splits that divide a region using that covariate. The mean of the variable importance across all trees yields the variable importance for the forest. For the SDForest, we report the reduction of the spectral loss (3) instead of the MSE. For the SDForest, the four most important covariates are also the four true causal parents of $Y$. Three of those also have a clear separation from the remaining 497 covariates. For the classical Random Forest, the variable importance for the true causal parents lies within the bulk of spurious covariates. The most important covariate among the true causal parents is only on rank 102, and we would have to use almost all the covariates to include all the causal parents in the selected set.
@@ 5.1 Qualitative Results
In addition to screening for the sparse causal set among a large number of covariates, we can also look at the functional dependence of the response $Y$ on the causal parents. We use partial dependence plots (Friedman 2001) to visualize the partial dependence. The idea is to predict $\hat{f}(\mathbf{X})$ for each observation and vary $\mathbf{X}_{j}$ over an interval, while keeping the other covariates at their observed values. This yields a different function of $X_{j}$ for every observation. The mean over all the observations can then be shown as a representative marginal effect of $X_{j}$ on the response $Y$.
Figure 5 shows how the SDForest estimates the true causal function. Especially for covariate 34, the estimated function closely approximates the true function. Covariate 108, the covariate with only slightly higher importance than the bulk of spurious covariates, has almost a constant influence on $Y$. This shows some limitations of estimating a sparse causal relationship in the presence of hidden confounding and high dimensionality. Estimating a nuanced true function, such as for variable 108, might be difficult due to the additional disturbances of confounding and noise. The classical Random Forest distributes the estimated function among all the spurious covariates and estimates almost a constant function for all four causal parents. The partial dependence plots for the estimated classical Random Forest are shown in Figure 6.
@@
Figure 6: Partial dependence plots as in Figure 5 but now with the estimated classical Random Forest by ranger for the four true causal parents of the response $Y$. The classical Random Forest essentially leads to constant functions, whereas the true dashed lines vary.
5.2 Quantitative Results
To quantify the performance of SDForests in estimating the causal function, we conduct simulation studies that vary different dimensions in the simulation. The default dimensions are $n=500$, $p=500$, and $q=20$, and we randomly draw data according to model (13). Each of those dimensions is varied separately to estimate the dependence of the performance on these different factors. Every experiment is repeated 200 times, where the entire data-generating process is redrawn at random each time. The performance is measured by the mean distance of the true causal function and the estimated function evaluated at 500 test observations $f_{mse}\coloneqq\frac{1}{n_{test}}\sum_{i=1}^{n_{test}}(f^{0}(x_{test,i})-\hat{f}(x_{test,i}))^{2}$.
The performance of the SDForests and the classical Random Forests depending on different dimensions is shown in Figure 7. Theorem 1 shows how the spectral objective (3) behaves as the number of variables $p$, and the number of observations $n$ grow to infinity. The term $R_{n}$ goes asymptotically with $1/\min(\sqrt{n},\sqrt{p})$ to zero. So, we need both $n$ and $p$ to grow for consistency. In this simulation study, we examine the dependence of performance on these quantities in practice. Figure 7a shows the error distribution depending on the sample size. The SDForests clearly outperforms the classical Random Forests in estimating the causal function, and the error decreases with a larger sample size.
We see the dependence of the performance on the number of covariates in Figure 7b. At $p>20$, we see how the SDForests start to perform significantly better than the classical Random Forests. When increasing $p$ above 50, the error no longer decreases, suggesting that there is a certain threshold of $p$ beyond which deconfounding is successful. Even with only the four causal parents as covariates, the SDForests do not lose performance in comparison to the classical Random Forests.
With stronger confounding, estimating the causal function becomes increasingly difficult. Not only does the bias increase, but the variance in the response also increases. Therefore, we expect the error to increase with increasing confounding even for the SDForests. Figure 7c shows this behavior with the simulation study, where we increase the number of confounding variables. For $q=0$, the setting corresponds to the classical model without confounding, and both the classical Random Forests and the SDForests perform equally well. It is important to note that we do not lose much, even when the data is not confounded, by applying the spectral deconfounding. However, we can gain a lot if there is confounding present.
In Figure 7d, we follow up on the assumption of dense confounding. Here, instead of having random effects of the confounders on all the covariates, we simulate the data with only a random subsample of the covariates being affected by confounding. The denser the confounding becomes, the better the SDForests perform. With around 200 covariates affected (40%), the SDForests start performing similarly as if there were no confounding, while with sparser confounding, SDForest still outperform the classical Random Forests.
@@ 6 Single-Cell Data
We apply SDForest and ranger to compare the resulting model on the scRNA-seq single-cell gene expression dataset for the cell RPE1 generated by Replogle et al. (2022). We use the preprocessed and filtered dataset provided by Chevalley et al. (2025). As the response variable, we choose the gene EIF1 following the arguments of Shen et al. (2023), who conjecture that EIF1 might be a leaf node in the causal graph of genes. We use all other gene expressions ($p=382$) as predictors. We consider observational data without any interventional gene knockouts ($n=11485$ observations).
Examining the singular values of the scaled predictors, as shown in Figure 8, reveals a clear spike in the first few singular values. This is an indication of a factor structure that may induce dense confounding.
Figure 8: Singular values of the scaled predictors in the single-cell data.
For both methods SDForest and ranger, we fit 500 trees and compare the estimated models in Figure 9. Both models are fitted with $\mathsf{mtry}=191$ and 1000 bootstrap samples per tree. The two methods agree on the variable importance for the most important genes. Thus, the most predictive genes, derived from ranger, are not just arising due to dense confounding. When comparing the predictions (or predicted regression functions), there is a substantial correlation between the methods, but a slight shrinkage towards zero is observed with SDForest due to optimizing the deconfounded loss, which shrinks the largest singular values. As a practical guideline, we advocate running and comparing SDForest and ranger, as it yields additional information and insights into whether variables and predictions are possibly induced by dense confounding or not. For this dataset, we do not have strong evidence for dense confounding.
Figure 9: Comparison of the classical Random Forest estimated by ranger and the SDForest in the single-cell data. Subfigure a) shows the variable importance for both methods scaled to the interval $[0,1]$. In subfigure b), the predictions $\hat{Y}$ of both methods are shown. The dashed 45-degree line corresponds to equality.
To test for robustness against dense confounding, we synthetically perturb the original cleaned and filtered dataset. For this, we construct $\mathbf{X}_{\tau}\coloneqq\mathbf{X}+\mathbf{H}\Gamma\tau$ and $\mathbf{Y}_{\tau}\coloneqq Y+\mathbf{H}\delta\tau$, where $\mathbf{Y}$ is the original expression of EIF1 and $\mathbf{X}$ are the other original gene expressions used for prediction. The entries of $\mathbf{H}\in\mathbb{R}^{n}$, $\Gamma\in\mathbb{R}^{1\times 382}$, and $\delta\in\mathbb{R}$ are sampled i.i.d. from $\mathcal{N}(0,1)$ while we vary $\tau$ to increase the added dense confounding. We fit both methods to $\mathbf{X}_{\tau}$ and $\mathbf{Y}_{\tau}$ as $\tau$ increases and analyze how the estimated functions change. The change in the estimated functions is calculated as the change of the out-of-bag predictions of $f^{\text{\emph{ranger}}}_{\tau}(\mathbf{X}^{\tau})$ and $f^{\text{SDF}}_{\tau}(\mathbf{X}^{\tau})$ depending on $\tau$
@@
We repeat the synthetic confounding of the data by sampling 20 times $\mathbf{H}$, $\Gamma$, and $\delta$ and increasing $\tau$. The distributions of the function changes are shown in Figure 10. With no added confounding, the difference between the two functions is small (see caption of Figure 10). This suggests that there is no major dense confounding present. As we increasingly add dense confounding, we clearly see how the estimated function by ranger changes, while SDForest demonstrates robustness against this added perturbation.
@@ 7 Conclusion
We propose the Spectrally Deconfounded Random Forest algorithm SDForest with R-package SDModels (Ulmer & Scheidegger 2025) to estimate direct regression functions in high-dimensional data in the presence of dense hidden confounding. This can be used to screen for relevant covariates among a large set of variables, and the procedure provides robustness against unobserved confounding, gaining much in the confounded case but losing little if no confounding is present. SDModels provides functions such as regularization paths and stability selection to screen for relevant covariates, as well as partial dependence plots to understand their relationship with the variable of interest. We demonstrate the empirical behavior of SDForests in various settings, including challenging cases with sparse confounding, to illustrate some fundamental limitations (which are not found to be clearly worse than those of classical Random Forests).
Many potentially confounded high-dimensional applications are about classification instead of regression. Currently, SDForest is only applicable for regression tasks, and it is important to extend this methodology to classification. Another open question is whether one can combine spectral deconfounding with Quantile Regression Forests (Meinshausen 2006) or Distributional Random Forests (Hothorn & Zeileis 2021, Ćevid et al. 2022) to gain access to prediction intervals or construct confidence intervals using techniques as in Guo et al. (2022) for the linear case and Näf et al. (2023) using Random Forests.
@@ Acknowledgment
We thank Mathieu Chevalley for his support in applying our method to the single-cell dataset. We would also like to thank Andrea Nava, Malte Londschien, Minhui Jiang, and Christoph Schultheiss for their helpful discussions and comments.
@@ Appendix A Approximation of splitting criteria
In the $M$th iteration of Algorithm 1, we need to find, in each region, the split that reduces the loss the most. This means we have $M$ regions in each iteration, each with an optimal split to choose from. When we split the region at the optimal point, we obtain two new regions. Now, due to the spectral transformation, the samples, as well as the different regions containing a subset of the samples, are not independent. Therefore, to truly find the next best split, we would need to determine the optimal split and its corresponding loss decrease in each region anew.
This is done in Figure 11 using the method SDT2. In our studies and as the default in the R-package, we only estimate the optimal split and its loss decrease for the two new regions and reuse all the other estimates from the previous iteration. This saves substantial computational time, and we argue that a previously good split stays reasonable. We compare the performance of the two options in Figure 11. We simulate data following the confounding model (1) and using the same parameters as in Section 5, but with a random regression tree for $f^{0}$ as in Equation 8. The random regression tree is grown using random splits until there are ten leaf nodes. While the reestimation of all splits in SDT2 might be a bit better, we do not see a significant decrease in performance when using the computationally much more efficient approximation SDT1.
In the case of smooth underlying regression functions, we expect to see similar behavior (because we can consider the approximation error of a smooth function with a tree that remains unaffected and the estimation error of the best-approximated tree function, which is analogous to the discussion above).